OpenStudy (anonymous):

Express the function F defined by F(x)=x^2/(x^2+3) in the form f of g for suitable f and g Help please

OpenStudy (irishboy123):

tried taking out x^2 as a separate function g(x) ?? or even x^2 + 3 ??

OpenStudy (anonymous):

let f(x) = x^2+3 and g(x) = x^2/x

OpenStudy (anonymous):

\[f(x)=x^2\] \[g(x)=\frac{x}{x+3}\]

OpenStudy (anonymous):

See if that works

OpenStudy (anonymous):

I was close, I said f(x) = x^2 and g(x) = \[\frac{ x }{ x ^{2}+3}\]